# Topological components of surface group representations and the Signature

### v1

@@ Introduction
In the theory of surface group representations, understanding the topology of representation spaces is a fundamental topic which influenced many branches of mathematics including non-abelian Hodge theory, Higgs bundle theory and many others. In particular, determining the number of connected components of these spaces is a basic yet crucial question. Topological invariants serve as powerful tools in addressing such problems. Among these invariants, the Toledo invariant plays a central role in the representation theory of surface groups [19]. A foundational result in this direction was established by Domic and Toledo [6], who proved the celebrated Milnor–Wood inequality. This landmark result was further generalized by Burger, Iozzi, and Wienhard [4], who extended the definition of the Toledo invariant to surfaces with boundary for any Hermitian Lie group, utilizing the framework of bounded cohomology. They also showed that the Milnor–Wood inequality continues to hold in this broader setting.
@@
In his seminal work [8], Goldman introduced the notion of the relative Euler class for boundary non-elliptic representations, showing it takes integer values. In our previous work [11], we proved that the Toledo invariant coincides precisely with the relative Euler class (see Proposition 1.2). By exploiting both the continuity of the relative Euler class and its integrality in the case of closed surfaces, Goldman [8, Theorem A] proved that the representation space $\mathrm{Hom}(\pi_{1}(\Sigma_{g,0}),\mathrm{SL}(2,\mathbb{R}))$ has exactly $2^{2g+1}+2g-3$ connected components. Goldman further extended these ideas to surfaces with boundary, showing that the space of representations in $\mathrm{PSL}(2,\mathbb{R})$ with hyperbolic boundary holonomy decomposes into connected components indexed by the relative Euler class [8, Theorem D]. However, much less is known about the topology of representation spaces with elliptic or parabolic boundary holonomies. In particular, determining the number of connected components in these cases remains a natural and open problem.
@@
In this paper, we address this problem by employing the signature invariant to systematically study the connected components of surface group representations into $\mathrm{SL}(2,\mathbb{R})$ and $\mathrm{PSL}(2,\mathbb{R})$ with fixed boundary holonomies of elliptic, hyperbolic, or parabolic type, and we aim to count the number of connected components of these representation spaces. A key difficulty arises from the fact that, for surfaces with boundary, the Toledo invariant is generally no longer an integer [4, Theorem 1(2)], making it insufficient to distinguish between connected components. Instead, we turn to the signature invariant as a more refined topological invariant that remains integer-valued and is well-suited for our purpose.
@@
In [9], Hitchin studied representations of closed surface groups into $G=\mathrm{SL}(2,\mathbb{R})$, and established a real-analytic correspondence between irreducible representations $\rho:\pi_{1}(\Sigma)\to\mathrm{SL}(2,\mathbb{C})$ and stable Higgs bundles. Since then, the Higgs bundle approach has become a powerful and widely adopted method for studying moduli spaces of surface group representations (see, for example, [5, 7, 17, 18]). In particular, for surfaces with punctures, the topology of the moduli space of Higgs bundles has been explored in the works such as [3, 15, 16]. However, these studies primarily focus on the case of nonzero relative Euler classes and do not compute the precise number of connected components (see e.g. [15]).
@@
where $\bm{\rho}(\phi)$ denotes the rho invariant (see Section 1.3), and $\mathrm{T}(\phi)$ is the Toledo invariant [10, Theorem 4]. If $\Sigma$ is the connected sum of two (connected) surfaces $\Sigma_{i}$ along a separating loop, then the signature satisfies
@@ 0.1. Boundary elliptic representations
be the natural projection. As shown in [8, Lemma 2.2], this induces a covering map onto its inverse image of each connected component:
@@ 0.2. Boundary hyperbolic representations
see [8, Theorem D]. However, the number of connected components of the corresponding space in $\mathrm{SL}(2,\mathbb{R})$, i.e., $\mathrm{Hyp}(\Sigma_{g,n})$, was not computed in that work. For any representation $\phi\in\mathrm{Hyp}(\Sigma_{g,n})$, its rho invariant vanishes. Consequently, the signature of $\phi$ satisfies
@@ 0.3. Boundary parabolic representations
$\{\mathrm{Par}(\Sigma_{g,n})\}=\begin{cases}4^{g+1}&\text{if }g+n=2,\\ 2^{2g+2n}+2^{2n-1}(4g+n-5+(n\bmod 2))&\text{otherwise}.\end{cases}$
@@
$\{\mathrm{Par}_{P}(\Sigma_{g,n})\}=\begin{cases}2^{g+1},&\text{if }g+n=2,\\ 2^{n}(4g+n-3+(n\bmod 2)),&\text{otherwise}.\end{cases}$
@@
Acknowledgments. Research by Inkang Kim is partially supported by Grant NRF-2019R1A2C1083865 and KIAS Individual Grant (MG031408). The first author is grateful for the support of IHES and MPI during his visits. Research by Xueyuan Wan is supported by Science and Technology Innovation Key R&D Program of Chongqing and the National Key R&D Program of China (Grant No. 2024YFA1013200) and the Natural Science Foundation of Chongqing, China (Grant No. CSTB2024NSCQ-LZX0040, CSTB2023NSCQ-LZX0042).
@@ 1.2.1. Toledo invariant
We now recall the definition of the Toledo invariant for surfaces with boundary, as introduced by Burger, Iozzi, and Wienhard in [4, Section 1.1]. The Toledo invariant can be expressed as the integral of the first Chern form with compact support over the surface (cf. [10]).
@@
According to [4, Theorem 1], for a group $G$ of Hermitian type and a representation $\phi:\pi_{1}(\Sigma)\to G$, the Milnor-Wood inequality holds:
@@ 1.3. Rho invariant
We now define the rho invariant associated to the group $\mathrm{SL}(2,\mathbb{R})$. The following definition is due to [10, Definition 4.1].
@@
By [10, Lemma 4.2, Lemma 4.3, and Corollary 4.4], the quantity $\rho(L,\mathbf{J},W)$ does not depend on the choices of $\mathbf{J}$ and $W$. Hence, we simply write:
@@
By [10, Theorem 1], for a compact oriented surface $\Sigma$ with nonempty boundary and a representation $\phi:\pi_{1}(\Sigma)\to\mathrm{SL}(2,\mathbb{R})$, the signature, Toledo invariant, and rho invariant satisfy the relation:
@@
see [10, Theorem 4].
Note that the Toledo invariant $\mathrm{T}(\phi)$ is continuous (see [4, Theorem 1]), and the rho invariant $\bm{\rho}(\phi)$ is continuous away from elements with eigenvalue 1 (cf. [10, Section 4]). Consequently, the signature $\mathrm{sign}(\phi)$ is locally constant and only jumps when the boundary holonomy acquires 1 as an eigenvalue.
@@ 2. Boundary elliptic representations
for some angle $\theta_{i}\in(0,\pi)\cup(\pi,2\pi)$ (resp. $\theta_{i}\in(0,2\pi)$), see [10, Definition 4.5].
@@ 2.1.1. Boundary elliptic-unipotent representations
Since the rho invariant is continuous on $\mathrm{SO}(2)\setminus\{I\}$ (see [10, Section 4]), the boundary rho invariant $\bm{\rho}(\phi_{t})$ varies continuously for any path $\{\phi_{t}\}_{t\in[0,1]}\subset\mathrm{E}(\Sigma)$. On the other hand, the Toledo invariant vanishes for every representation $\phi\in\mathrm{Hom}(\pi_{1}(\Sigma),\mathrm{SO}(2))$ (see [10, Section 6.2]). Therefore, for all such paths, we have
@@
Note that when $n=1$, $\mathrm{E}(\Sigma)=\emptyset$. Hence, we assume $n\geq 2$ in the following. It is shown in [11, Theorem 2] that
@@ 2.2.1. Boundary elliptic-unipotent representations
Therefore, $\mathrm{Ell}(\Sigma)$ is a dense subset of $\mathrm{E}(\Sigma)$. In particular, any boundary elliptic-unipotent representation can be deformed into a boundary elliptic representation by a small perturbation, while the signature remains unchanged under such deformation. By [11, Theorem 4], we have
@@
If $|\mathrm{tr}(C_{1}\cdots C_{n-2})|<2$, then the product is elliptic, and we can deform $\phi$ within $\mathrm{Ell}(\Sigma)$ while keeping eigenvalues fixed, such that $C_{1}\cdots C_{n-2}\in\mathrm{SO}(2)\setminus\{\pm I\}$. By Lemma 2.6, we can fix $C_{1},\ldots,C_{n-2}$ and deform $C_{n-1}$ into $\mathrm{SO}(2)\setminus\{\pm I\}$ through a path $C_{n-1}(t)$ satisfying $|\mathrm{tr}(C_{1}\cdots C_{n-2}C_{n-1}(t))|<2$. Then $C_{n}(t):=(C_{1}\cdots C_{n-2}C_{n-1}(t))^{-1}\in\mathrm{SO}(2)\setminus\{\pm I\}$, and the resulting deformed representation $\phi^{\prime}$ satisfies $\phi^{\prime}(c_{n-1}),\phi^{\prime}(c_{n})\in\mathrm{SO}(2)\setminus\{\pm I\}$.
@@
Thus, after a slight perturbation and a reparameterization, we may deform $\phi$ to a new representation $\phi^{\prime}$ such that $\phi^{\prime}(c_{n-1})=C_{n-1}(1)\in\mathrm{SO}(2)\setminus\{\pm I\}$, or $C_{n-1}(1)$ is sufficiently close to $\pm I$ so that $|\mathrm{tr}(C_{n-2}C_{n-1}(1))|<2$. We also ensure that $\phi^{\prime}(c_{1})=\phi(c_{1})\in\mathrm{SO}(2)\setminus\{\pm I\}$.
@@ 2.2.2. Boundary elliptic representations
Similarly, one may consider boundary elliptic representations. According to [11, Theorem 4], we have
@@ 2.3.2. The case of $n=2$
In this section, we assume that $g=1$ and $n=2$. In this case, by [11, Theorem 4], the possible values of the signature for boundary elliptic (or boundary elliptic-unipotent) representations are given by the set $\{-4,0,4\}$.
@@
By a slight perturbation, we may assume that $|\mathrm{tr}([A,B])|\neq 2$. If $\mathrm{tr}([A,B])>2$, then by the proof of Lemma 2.12, the commutator $[A,B]$ can be deformed to the identity, which implies $\mathrm{sign}(\phi_{1})=0$, a contradiction. See also [11, Proposition 9.7] for further discussion.
@@ 2.4. The case of general genus
for all distinct indices $1\leq i_{1},\dots,i_{r}\leq n$. Up to conjugation, we may assume $C_{1}\in\mathrm{SO}(2)\setminus\{\pm I\}$.
@@
If $C_{0}$ is hyperbolic, we glue the surface $\Sigma_{g,1}$ with a surface $\Sigma_{1,1}$ along $c_{0}$ to obtain a closed surface $\Sigma_{g+1,0}$. Let $a^{\prime},b^{\prime},c_{0}$ be generators of $\pi_{1}(\Sigma_{1,1})$. Extending $\phi_{1}$ naturally defines a representation $\widetilde{\phi_{1}}$ on $\pi_{1}(\Sigma_{g+1,0})$. By [8, Proof of Theorem A], the representation in $\mathrm{Hom}(\pi_{1}(\Sigma_{g+1,0}),\mathrm{SL}(2,\mathbb{R}))$ with maximal or minimal signature (i.e., $\mathrm{sign}=\pm 2|\chi(\Sigma_{g+1,0})|$) can be deformed to a representation $\widetilde{\psi}$ such that $\widetilde{\psi}([a^{\prime},b^{\prime}])\in\mathrm{SO}(2)\setminus\{\pm I\}$. Furthermore, all representations with neither maximal nor minimal signature lie in a single connected component, and can therefore be deformed to a representation $\widetilde{\psi}$ with $\widetilde{\psi}([a^{\prime},b^{\prime}])=I$. In one word, the representation $\widetilde{\phi\\ _{1}}$ can be deformed to some $\widetilde{\psi}$ satisfying $\widetilde{\psi}([a^{\prime},b^{\prime}])\in\mathrm{SO}(2)$. Restricting $\widetilde{\psi}$ to $\pi_{1}(\Sigma_{g,1})$ yields a deformation $\phi_{1}(t)$ of $\phi_{1}$ such that $\phi_{1}(t)(c_{0})\in\mathrm{SO}(2)$.
Let $P(t):=\phi_{1}(t)(c_{0})$, which lies in $\mathrm{SO}(2)$ at $t=1$. We now construct deformations $C_{1}(t),\dots,C_{n}(t)$ such that
(2.14) $C_{1}(t)\cdots C_{n}(t)=P(t),\quad\text{and}\quad C_{i}(1)\in\mathrm{SO}(2)\setminus\{\pm I\}.$
For $n=2$, the existence of such $C_{i}(t)$ with conjugacy class of $C_{1}$ fixed follows from Lemma 2.12. Assume the claim holds for n1n-1, and that the conjugacy class of one $C_{i}$ can be fixed during deformation.
If $C_{2}C_{3}$ is hyperbolic, then by Proposition 2.7, we can deform $C_{2}(t)$ and $C_{3}(t)$ such that $C_{2}(1)\in\mathrm{SO}(2)\setminus\{\pm I\}$ or very closed to $\pm I$, and $C_{2}(t)C_{3}(t)$ is constant. We may then regard $C_{2}^{\prime}:=C_{1}C_{2}$ as elliptic. Thus, equation (2.14) reduces to the (n1)(n-1)-term case, and we can find deformations $C_{2}^{\prime}(t),C_{3}(t),\dots,C_{n}(t)$ such that $C_{2}^{\prime}(1),C_{3}(1),\dots,C_{n}(1)\in\mathrm{SO}(2)\setminus\{\pm I\}$ and the conjugacy class of $C_{2}^{\prime}$ is preserved. By Lemma 2.6, we can fix $C_{2}^{\prime}$ and the conjugacy class of $C_{1}$, and obtain deformation $C_{1}^{\prime\prime}(t)$ (resp. $C_{2}^{\prime\prime}(t)$)of $C_{1}$ (resp. $C_{2}$) such that $C_{2}^{\prime}=C_{1}^{\prime\prime}(t)C_{2}^{\prime\prime}(t),$ where $C_{2}^{\prime}$, $C_{1}^{\prime\prime}(1)$, and $C_{2}^{\prime\prime}(1)$ commute. Suppose $C_{2}^{\prime}(t)=U(t)C_{2}^{\prime}U(t)^{-1}$; then define
@@
With this construction, the conjugacy class of $C_{1}(t)$ is preserved, and we have $C_{1}(t)C_{2}(t)=C_{2}^{\prime}(t).$ Moreover, since $C_{2}^{\prime}(1)\in\mathrm{SO}(2)\setminus\{\pm I\}$ and the matrices $C_{1}^{\prime\prime}(1),C_{2}^{\prime\prime}(1),C_{2}^{\prime}(1)$ commute, it follows that $C_{1}^{\prime\prime}(1),C_{2}^{\prime\prime}(1)\in\mathrm{SO}(2)\setminus\{\pm I\}.$
If $C_{2}C_{3}$ is elliptic, we treat it as a single unit, thereby reducing the situation to the previous case. The proof is complete. ∎

### v3

# Topological components of surface group representations and signature

@@ Introduction
In the theory of surface group representations, understanding the topology of representation spaces is a fundamental topic which influenced many branches of mathematics including non-abelian Hodge theory, Higgs bundle theory and many others [11, 26, 6]. In particular, determining the number of connected components of these spaces is a basic yet crucial question [10, 12]. Topological invariants serve as powerful tools in addressing such problems. Among these invariants, the Toledo invariant plays a central role in the representation theory of surface groups [27, 5, 4]. A foundational result in this direction was established by Domic and Toledo [8], who proved the celebrated Milnor–Wood inequality [20, 28]. This landmark result was further generalized by Burger, Iozzi, and Wienhard [5], who extended the definition of the Toledo invariant to surfaces with boundary for any Hermitian Lie group, utilizing the framework of bounded cohomology. They also showed that the Milnor–Wood inequality continues to hold in this broader setting.
@@
In his seminal work [10], Goldman introduced the notion of the relative Euler class for boundary non-elliptic representations, showing it takes integer values. In our previous work [15], we proved that the Toledo invariant coincides precisely with the relative Euler class (see Proposition 1.2). By exploiting both the continuity of the relative Euler class and its integrality in the case of closed surfaces, Goldman [10, Theorem A] proved that the representation space $\mathrm{Hom}(\pi_{1}(\Sigma_{g,0}),\mathrm{SL}(2,\mathbb{R}))$ has exactly $2^{2g+1}+2g-3$ connected components. Goldman further extended these ideas to surfaces with boundary, showing that the space of representations in $\mathrm{PSL}(2,\mathbb{R})$ with hyperbolic boundary holonomy decomposes into connected components indexed by the relative Euler class [10, Theorem D]. Recently, Ryu and Yang [23] investigated the topology of the connected components of boundary parabolic representations into $\mathrm{PSL}(2,\mathbb{R})$ and determined their number, see also [7, 13, 17, 29] for related discussions. In contrast, much less is known about the topology of representation spaces with elliptic boundary holonomies. In particular, determining the number of connected components in this setting remains open.
@@
In this paper, we address this problem by employing the signature invariant to systematically study the connected components of surface group representations into $\mathrm{SL}(2,\mathbb{R})$ and $\mathrm{PSL}(2,\mathbb{R})$ with fixed boundary holonomies of elliptic, hyperbolic, or parabolic type, and we aim to count the number of connected components of these representation spaces. A key difficulty arises from the fact that, for surfaces with boundary, the Toledo invariant is generally no longer an integer [5, Theorem 1(2)], making it insufficient to distinguish between connected components. Instead, we turn to the signature invariant as a more refined topological invariant that remains integer-valued and is well-suited for our purpose.
@@
In [11], Hitchin studied representations of closed surface groups into $G=\mathrm{SL}(2,\mathbb{R})$, and established a real-analytic correspondence between irreducible representations $\rho:\pi_{1}(\Sigma)\to\mathrm{SL}(2,\mathbb{C})$ and stable Higgs bundles. Since then, the Higgs bundle approach has become a powerful and widely adopted method for studying moduli spaces of surface group representations (see, for example, [6, 9, 24, 25]). In particular, for surfaces with punctures, the topology of the moduli space of Higgs bundles has been explored in works such as [3, 21, 22]. However, these studies primarily focus on the case of nonzero relative Euler classes and do not compute the precise number of connected components (see e.g. [21]).
@@
where $\bm{\rho}(\phi)$ denotes the rho invariant (see Section 1.3), and $\mathrm{T}(\phi)$ is the Toledo invariant [14, Theorem 4]. If $\Sigma$ is the connected sum of two (connected) surfaces $\Sigma_{i}$ along a separating loop, then the signature satisfies
@@ 0.1. Boundary elliptic representations
be the natural projection. As shown in [10, Lemma 2.2], this induces a covering map onto its inverse image of each connected component:
@@ 0.2. Boundary hyperbolic representations
see [10, Theorem D]. However, the number of connected components of the corresponding space in $\mathrm{SL}(2,\mathbb{R})$, i.e., $\mathrm{Hyp}(\Sigma_{g,n})$, was not computed in that work. For any representation $\phi\in\mathrm{Hyp}(\Sigma_{g,n})$, its rho invariant vanishes. Consequently, the signature of $\phi$ satisfies
@@ 0.3. Boundary parabolic representations
[…]
@@
$\{\mathrm{Par}_{P}(\Sigma_{g,n})\}=\begin{cases}2^{n}(4g+n-3)&\text{for }g\geq 1\\ 2^{n}\big{(}n-3\big{)}+2(n+1)&\text{for }g=0.\end{cases}$
Remark 0.6.
In [23, Corollary 1.5], Ryu and Yang proved Theorem 0.5 by a completely different method. Their approach was to compute the number of connected components corresponding to each relative Euler class, and then sum these numbers to obtain the total.
Our method instead proceeds by deforming to interior hyperbolic representations, thereby establishing a correspondence between boundary parabolic and boundary hyperbolic representations. Using the counting formula for boundary hyperbolic representations, we then deduce the number of connected components.
Here we briefly discuss some potential future research topics related to the study of topological components of surface group representations.
Let $\Sigma=\Sigma_{g,n}$ be a general surface with boundary. One may consider a surface group representation $\phi:\pi_{1}(\Sigma)\to G$, where $G$ is a $k$-fold covering $\pi:G\rightarrow\mathrm{PSL}(2,\mathbb{R})$. A natural problem is to study the topological components of the spaces $\mathrm{Hom}(\pi_{1}(\Sigma),G)$ consisting of boundary elliptic, hyperbolic, and parabolic representations, and to compute the number of connected components. When $\Sigma$ is closed, i.e., $n=0$, Goldman [10, Theorem A] has obtained the count of connected components in this case. For surfaces with boundary, this paper mainly considers the cases $k=1,2$. Clearly, when $g=0$, one can easily obtain a counting formula since $\pi_{*}:\mathrm{Hom}(\pi_{1}(\Sigma_{0,n}),G)\to\mathrm{Hom}(\pi_{1}(\Sigma_{0,n}),\mathrm{PSL}(2,\mathbb{R}))$ is a covering map of degree $k^{n-1}$. However, for general $g\geq 1$, the situation is more complicated compared with these two extreme cases, arising from the action $\mathrm{Hom}(\pi_{1}(\Sigma_{g,n}),\mathrm{Ker}(\pi))$.
In [14], together with Pansu, we established a signature formula for surface group representations into $\mathrm{U}(p,q)$, and obtained a Milnor–Wood type inequality for the signature. Moreover, we introduced the definitions of boundary elliptic-unipotent, hyperbolic-unipotent, and unipotent (or parabolic) representations. Thus, a natural problem is to study the topological properties and to count the connected components of these subspaces. Note that $\mathrm{SL}(2,\mathbb{R})$ is isomorphic to $\mathrm{SU}(1,1)$, so from the results of this paper, the case $p=q=1$ is readily understood. However, for general $(p,q)$, this remains an open and interesting question. In addition, when $q=0$, i.e., in the $\mathrm{U}(p)$ case, the Toledo invariant is identically zero. In [15], we completely determined the possible values of signature, which makes it natural to investigate the topological components of boundary elliptic representations into $\mathrm{U}(p)$. Note that when $p=1$, $\mathrm{U}(1)$ is isomorphic to $\mathrm{SO}(2)$, which is studied systematically in Section 2.1, and Goldman [10, Theorem A (ii)] proved that for closed surfaces, the space of surface group representations into $\mathrm{SU}(2)$ is connected.
@@
Acknowledgments. Research by Inkang Kim is partially supported by Grant NRF-2019R1A2C1083865 and KIAS Individual Grant (MG031408). The first author is grateful for the support of IHES and MPI during his visits. Research by Xueyuan Wan is supported by Science and Technology Innovation Key R&D Program of Chongqing and the National Key R&D Program of China (Grant No. 2024YFA1013200) and the Natural Science Foundation of Chongqing, China (Grant No. CSTB2024NSCQ-LZX0040, CSTB2023NSCQ-LZX0042). The second author would like to thank Qiongling Li for bringing [23] to his attention.
@@ 1.2.1. Toledo invariant
We now recall the definition of the Toledo invariant for surfaces with boundary, as introduced by Burger, Iozzi, and Wienhard in [5, Section 1.1]. The Toledo invariant can be expressed as the integral of the first Chern form with compact support over the surface (cf. [14]).
@@
According to [5, Theorem 1], for a group $G$ of Hermitian type and a representation $\phi:\pi_{1}(\Sigma)\to G$, the Milnor-Wood inequality holds:
@@ 1.3. Rho invariant
We now define the rho invariant associated to the group $\mathrm{SL}(2,\mathbb{R})$. The following definition is due to [14, Definition 4.1].
@@
By [14, Lemma 4.2, Lemma 4.3, and Corollary 4.4], the quantity $\rho(L,\mathbf{J},W)$ does not depend on the choices of $\mathbf{J}$ and $W$. Hence, we simply write:
@@
By [14, Theorem 1], for a compact oriented surface $\Sigma$ with nonempty boundary and a representation $\phi:\pi_{1}(\Sigma)\to\mathrm{SL}(2,\mathbb{R})$, the signature, Toledo invariant, and rho invariant satisfy the relation:
@@
see [14, Theorem 4].
Note that the Toledo invariant $\mathrm{T}(\phi)$ is continuous (see [5, Theorem 1]), and the rho invariant $\bm{\rho}(\phi)$ is continuous away from elements with eigenvalue 1 (cf. [14, Section 4]). Consequently, the signature $\mathrm{sign}(\phi)$ is locally constant and only jumps when the boundary holonomy acquires 1 as an eigenvalue.
@@ 2. Boundary elliptic representations
for some angle $\theta_{i}\in(0,\pi)\cup(\pi,2\pi)$ (resp. $\theta_{i}\in(0,2\pi)$), see [14, Definition 4.5].
@@ 2.1.1. Boundary elliptic-unipotent representations
Since the rho invariant is continuous on $\mathrm{SO}(2)\setminus\{I\}$ (see [14, Section 4]), the boundary rho invariant $\bm{\rho}(\phi_{t})$ varies continuously for any path $\{\phi_{t}\}_{t\in[0,1]}\subset\mathrm{E}(\Sigma)$. On the other hand, the Toledo invariant vanishes for every representation $\phi\in\mathrm{Hom}(\pi_{1}(\Sigma),\mathrm{SO}(2))$ (see [14, Section 6.2]). Therefore, for all such paths, we have
@@
Note that when $n=1$, $\mathrm{E}(\Sigma)=\emptyset$. Hence, we assume $n\geq 2$ in the following. A simple calculation shows that
@@ 2.2.1. Boundary elliptic-unipotent representations
Therefore, $\mathrm{Ell}(\Sigma)$ is a dense subset of $\mathrm{E}(\Sigma)$. In particular, any boundary elliptic-unipotent representation can be deformed into a boundary elliptic representation by a small perturbation, while the signature remains unchanged under such deformation. By [15, Theorem 4], we have
@@
If $|\mathrm{tr}(C_{1}\cdots C_{n-2})|<2$, then the product is elliptic, and we can deform $\phi$ within $\mathrm{Ell}(\Sigma)$ while keeping the conjugacy classes of $C_{1},\cdots,C_{n-2}$ fixed, such that the product $C_{1}\cdots C_{n-2}$ in $\mathrm{SO}(2)\setminus\{\pm I\}$. By Lemma 2.6, we can fix $C_{1},\ldots,C_{n-2}$ and deform $C_{n-1}$ into $\mathrm{SO}(2)\setminus\{\pm I\}$ through a path $C_{n-1}(t)$ satisfying $|\mathrm{tr}(C_{1}\cdots C_{n-2}C_{n-1}(t))|<2$. Then $C_{n}(t):=(C_{1}\cdots C_{n-2}C_{n-1}(t))^{-1}$satisfying $C_{n}(1)\in\mathrm{SO}(2)\setminus\{\pm I\}$, and the resulting deformed representation $\phi^{\prime}$ satisfies $\phi^{\prime}(c_{n-1}),\phi^{\prime}(c_{n})\in\mathrm{SO}(2)\setminus\{\pm I\}$.
@@
Thus, after a slight perturbation and a reparameterization, we may deform $\phi$ to a new representation $\phi^{\prime}$ such that $\phi^{\prime}(c_{n-1})=C_{n-1}(1)\in\mathrm{SO}(2)\setminus\{\pm I\}$, or $C_{n-1}(1)$ is sufficiently close to $\pm I$ so that $|\mathrm{tr}(C_{n-2}C_{n-1}(1))|<2$. We also ensure that $\phi^{\prime}(c_{n-2})=\phi(c_{n-2})\in\mathrm{SO}(2)\setminus\{\pm I\}$.
@@ 2.2.2. Boundary elliptic representations
Similarly, one may consider boundary elliptic representations. According to [15, Theorem 4], we have
@@ 2.3.2. The case of $n=2$
In this section, we assume that $g=1$ and $n=2$. In this case, by [15, Theorem 4], the possible values of the signature for boundary elliptic (or boundary elliptic-unipotent) representations are given by the set $\{-4,0,4\}$.
@@
By a slight perturbation, we may assume that $|\mathrm{tr}([A,B])|\neq 2$. If $\mathrm{tr}([A,B])>2$, then by the proof of Lemma 2.12, the commutator $[A,B]$ can be deformed to the identity, which implies $\mathrm{sign}(\phi_{1})=0$, a contradiction. See also [15, Proposition 9.7] for further discussion.
@@ 2.4. The case of general genus
for all distinct indices $1\leq i_{1},\dots,i_{r}\leq n$.
@@
If $C_{0}$ is hyperbolic, then since the space $\mathrm{Hom}(\pi_{1}(\Sigma_{g,1}),\mathrm{SL}(2,\mathbb{R}))$ is connected, there exists a path $\phi_{1}(t)$ starting from $\phi_{1}$ such that $\phi_{1}(1)(c_{0})\in\mathrm{SO}(2)$. Moreover, the path $\phi_{1}(t)$ can be chosen so that there exists a unique $t\in[0,1]$ at which $|\mathrm{tr}(\phi_{1}(t)(c_{0}))|=2$, and $\phi_{1}(1)(c_{0})=I$ if $\mathrm{tr}(C_{0})>2,$ $\phi_{1}(1)(c_{0})\in\mathrm{SO}(2)\setminus\{\pm I\}$ if $\mathrm{tr}(C_{0})<-2.$
Set $P(t):=\phi_{1}(t)(c_{0})$. We now construct deformations $C_{1}(t),\dots,C_{n}(t)$ such that
(2.14) $C_{1}(t)\cdots C_{n}(t)=P(t)^{-1},\quad\text{and}\quad C_{i}(1)\in\mathrm{SO}(2)\setminus\{\pm I\}.$
For $n=2$, the existence of such $C_{i}(t)$ with the conjugacy class of $C_{1}$ or $C_{2}$ fixed follows from the proof of Lemma 2.12. Assume that (2.14) holds for n1n-1, and that the conjugacy class of any one of the $C_{i}$ can be fixed during deformation. We now prove the case for $n$, under the condition that the conjugacy class of an arbitrarily chosen $C_{i}$ is fixed. Without loss of generality, we fix the conjugacy class of $C_{1}$ and solve (2.14) accordingly.
If $C_{2}C_{3}$ is hyperbolic, then by the proof of Proposition 2.7 ($C_{2},C_{3}$ play the role of $C_{n-1},C_{n}$ there), we can obtain the deformtion $\widetilde{C}_{i}(t)$ of $C_{i}$ for $i=2,3$ such that $\widetilde{C}_{2}(1)$ commutes with $C_{1}$ or very closed to $\pm I$, and $\widetilde{C}_{2}(t)\widetilde{C}_{3}(t)$ is constant. Up to a small perturbation on $\widetilde{C}_{2}(1)$, we may then regard $C_{2}^{\prime}:=C_{1}C_{2}$ as elliptic (when $\widetilde{C}_{2}(1)$ commutes with $C_{1}$, it is obvious, and when it is very close to $\pm I$, $|\mathrm{tr}(C_{1}\widetilde{C}_{2}(1))|<2$). Thus, equation (2.14) reduces to the (n1)(n-1)-term case, and we can find deformations $C_{2}^{\prime}(t),C_{3}(t),\dots,C_{n}(t)$ such that $C_{2}^{\prime}(1),C_{3}(1),\dots,C_{n}(1)\in\mathrm{SO}(2)\setminus\{\pm I\}$ and the conjugacy class of $C_{2}^{\prime}$ is preserved.
By Lemma 2.6 (using $C_{1}^{-1}C_{2}^{\prime}=C_{2}$ we obtain $C_{1}^{\prime\prime}(t)$ such that $C_{1}^{\prime\prime}(t)^{-1}C_{2}^{\prime}=C_{2}^{\prime\prime}(t)$), we can fix $C_{2}^{\prime}$ and the conjugacy class of $C_{1}$, and obtain deformation $C_{1}^{\prime\prime}(t)$ (resp. $C_{2}^{\prime\prime}(t)$) of $C_{1}$ (resp. $C_{2}$) such that $C_{2}^{\prime}=C_{1}^{\prime\prime}(t)C_{2}^{\prime\prime}(t),$ where $C_{2}^{\prime}$, $C_{1}^{\prime\prime}(1)$, and $C_{2}^{\prime\prime}(1)$ commute. Since the conjugacy class of $C_{2}^{\prime}(t)$ is invariant, there exists $U(t)\in\mathrm{SL}(2,\mathbb{R})$ with $U(0)=I$ such that $C_{2}^{\prime}(t)=U(t)C_{2}^{\prime}U(t)^{-1}$, then define
@@
With this construction, the conjugacy class of $C_{1}(t)$ is preserved, and we have $C_{1}(t)C_{2}(t)=C_{2}^{\prime}(t).$ Moreover, since $C_{2}^{\prime}(1)\in\mathrm{SO}(2)\setminus\{\pm I\}$ and the matrices $C_{1}(1),C_{2}(1),C_{2}^{\prime}(1)$ are commute, it follows that $C_{1}(1),C_{2}(1)\in\mathrm{SO}(2)\setminus\{\pm I\}.$ In one word, $C_{1}(t),C_{2}(t),C_{3}(t),\cdots,C_{n}(t)$ solve (2.14) and the conjucay class of $C_{1}(t)$ is preserved.
If $C_{2}C_{3}$ is elliptic, we treat it as a single unit, thereby reducing the situation to the previous case. More precisely, by Lemma 2.6 (again using $C_{2}^{-1}C_{2}C_{3}=C_{3}$), we may keep $C_{2}C_{3}$ fixed while deforming $C_{2}$ and $C_{3}$ so that, in the end, they commute. Hence, without loss of generality, we may assume that $C_{2}$ and $C_{3}$ commute. Suppose $C_{i}\sim R(\theta_{i})$ for $i=2,3$. Then we can deform $\theta_{2}$ and $\theta_{3}$ in such a way that $\theta_{2}+\theta_{3}$ remains invariant, while $R(\theta_{2})$ approaches $\pm I$. Consequently, the product $C_{1}C_{2}$ can be deformed to an elliptic element $C_{2}^{\prime}$. The remainder of the proof then follows exactly as in the previous case. Now the proof is complete. ∎